Write the equation of a line that passes through the point (-2, 1) and has a slope of 4
Answer by Cromlix(4381)

Using equation -> y - b = m(x - a)
y - 1 = 4(x -(-2)
y - 1 = 4(x + 2)
y -1 = 4x + 8
y = 4x + 8 + 1
y = 4x + 9
Hope this helps.
:-)